A wheel is initially at rest is subjected to an angular acceleration and it completes 10 rotations in time 't', then the time taken by it to complete next 10 rotations

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Ram Sai Charan,

Ans:- Let the angular acceleration is a

θ=1/2 a t²

or 10×2∏=1/2 a t²

a=40∏/t².......................(1)

let the total time taken by it to rotate 20 rotations=T Hence from 1 we get,

T²=80∏/a

=2t²

T=√2 t

Hence if we calculate the time from the instant it finished 10 rotations is =(√2 - 1)t

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Ram Sai Charan !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur